OpenStudy (vicky123):

solve (5/x+5)= (15/x+7)

OpenStudy (anonymous):

-5

OpenStudy (vicky123):

how did you get that?

OpenStudy (jack1):

\[ (5/x+5)= (15/x+7)\] so as they're both denominators, times each side by the denominator to bring it to the top: \[(5/x+5) \times (x+5) = (15/x+7) \times (x+5)\] \[ \frac{ 5(x+5) }{ x+5 } = (\frac{ 15(x+5) }{ x+7 })\]

OpenStudy (jack1):

so the x+5/x+5 on the LHS of the brackets cancel each other out by becoming 1/1 so you're left with 5 on the LHS then do the same thing with the x+7 multiply both sides of the equals sign by x+7

OpenStudy (anonymous):

search of x?